A car dealer advertised a big sale by stretching a string of banners from the top of the building to the edge of the driveway. If the building is 26 m high and the driveway is 43 m from the building, how long is the string of banners?
Answer by nerdybill(7384)

A car dealer advertised a big sale by stretching a string of banners from the top of the building to the edge of the driveway. If the building is 26 m high and the driveway is 43 m from the building, how long is the string of banners?
.
Apply Pythagorean theorem:
Let h = length of banner (the hypotenuse)
26^2 + 43^2 = h^2
676 + 1849 = h^2
2525 = h^2
50.25 meters = h (length of banners)